OpenStudy (anonymous):

The sum of two numbers is 60. Their difference is 12, what is the smaller number? 24 12 30 -12

OpenStudy (anonymous):

60-24=36 36-24=12

OpenStudy (mayankdevnani):

Let the two numbers be x and y and we don't know which number is greater or smaller.

OpenStudy (mayankdevnani):

so, \[\large \bf x+y=60\] \[\large \bf x-y=12\]

OpenStudy (mayankdevnani):

solve this SIMULTANEOUS EQUATIONS and get your answer :) @Daniella33

OpenStudy (mayankdevnani):

SORRY , x>y

OpenStudy (mayankdevnani):

can you solve this equations ? @Daniella33

OpenStudy (anonymous):

marry me :DDD

OpenStudy (anonymous):

Im still confused

OpenStudy (anonymous):

the answer's 24

OpenStudy (mayankdevnani):

lolzz.. @sourwing